OpenStudy (anonymous):

Write the following expression in terms of the first power of the sine only. sin z + cos z I am also having trouble with this one. Please help!

OpenStudy (anonymous):

y=sqrt(2)sin(x+pi/4)

OpenStudy (anonymous):

How did you solve this?

OpenStudy (anonymous):

replaces cosz with sqrt(1-sin^2(z))

OpenStudy (anonymous):

then add the two sins together and get 2 sin x? how did you know to add π/4?

OpenStudy (anonymous):

sorry wrong approach to do tht, lemme start over

OpenStudy (anonymous):

sin(90-x)=cosx

OpenStudy (anonymous):

sin(pi/2-x)=cosx

OpenStudy (anonymous):

im so confused! i dont understand how to write it in terms of sine

OpenStudy (anonymous):

um do you kno sin (a+b) , sin(a-b) formula? if u do then write cosz as sin(pi/2-z) and use the identity

OpenStudy (anonymous):

oh wait i just looped u lol

OpenStudy (anonymous):

lol. im trying to figure it out but its not going so well.

OpenStudy (anonymous):

i just woke up im sorrygive me two minutes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

k i want u to imagine tht sinx+cosx is written as sinxcos(pi/4)+cosxsin(pi/4) do u kno why chose pi/4, because sin(pi/4) and cos(pi/4) both =1. now we can rewrite sinxcos(pi/4)+cosxsin(pi/4) as sin(pi/4+x)

OpenStudy (anonymous):

i mean they sin(pi/4)=sqrt 2 /2 and same for cos god im sleepy lol

OpenStudy (anonymous):

\[\sin(x)+\cos(x)=>\sin(x)*\cos(\pi/4)+\cos(x)\sin(\pi/4)\]

OpenStudy (anonymous):

if im writing it is terms of sine then why would cosine still be there? that where im confused

OpenStudy (anonymous):

sin(x)∗cos(π/4)+cos(x)sin(π/4)=1/sqrt(2)(sinx+cosx)

OpenStudy (anonymous):

but it can also be written as sin(x+pi/4)=1/sqrt(2) (sinx+cosx)

OpenStudy (anonymous):

bring sqrt(2) to the left hand side, by multiplying both sides sqrt2, and u get the answer =)

OpenStudy (anonymous):

lemme kno what part u didnt get and ill reexplain

OpenStudy (anonymous):

so the answer is sqrt(2) sin(x+π/4) = sins + cosx

OpenStudy (anonymous):

does the sqrt(2) stay multiplied outside or do i multiply it to everything??

OpenStudy (anonymous):

it stays outside

OpenStudy (anonymous):

\[\sqrt(2) \sin(\pi/4 +x)\]

OpenStudy (anonymous):

goodluck

OpenStudy (anonymous):

thanks for your help!

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

not to be greedy or anything =P

OpenStudy (anonymous):

ill give you one! but whats the answer? im still confused whats the work?

OpenStudy (anonymous):

lol? i showed u the work